OpenStudy (anonymous):

during an intramural basketball game, Team A scored 14 fewer points than Team B. Together,both teams scored a total of 148 points. how many points did Team A score during the game.

OpenStudy (anonymous):

A=x-14 B=x x-14+x=148 2x=162 x=81 A therefore scored 81-14=67points

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

any time